Write the slope-intercept equation for the line that passes through (-5, 10) and (-4, 3). Please show all of your work.
Answer by stanbon(75887)

Write the slope-intercept equation for the line that passes through (-5, 10) and (-4, 3).
----
Form: y = mx + b
solve for "m" and "b":
m = (10-3)/(-5--4) = 7/-1 = -7
----
Substitute for y,x,m and solve for "b":
3 = -7*-4+b
3 = 28+b
b = -25
-----
Equation:
y = -7x-25
===============
Cheers,
Stan H.
===============